OpenStudy (sand-lock53):

Will medal and fan! 1.) which choice could be the equation of a line perpendicular to the line represented by this equation? y=2/5x+2 A.) y= 5/2x+7 B.) y= -5/2x-3 2.) which chioce could be the equation of a line parallel to the line represented by this equation? 7x-12y=-32 A.) y=7/12x-9 B.) y=7x-12 3.) wich choice is the equation of a line that has a y-intercept of -3 and is perpendicular to the line represented by this equation? y=2/3x+1 A.) y= 3x-3 B.) y= -3/2x-3 C.) y- 3/2x+3

OpenStudy (sand-lock53):

Last, but not least, 4.) which choice is the equation of a line that passes through (9,-3) and is parallel to the line represented by this equation? y=5/3x-4 A.) y=5/3x-18 B.) y=-3/5x-5 C.) y=5x-4

OpenStudy (rayep12):

The first one is B the second one is A

OpenStudy (rayep12):

The third on is b

OpenStudy (rayep12):

the last one is C